OpenStudy (anonymous):

How can I write 8a^9/125a^6 with only one exponent?

OpenStudy (anonymous):

Well, for starters, this holds true... \[\huge \frac{a^m}{a^n}=a^{m-n}\]:D

OpenStudy (anonymous):

divide it out

OpenStudy (anonymous):

So that means... \[\huge \frac{8a^9}{125a^6}=\frac{8a^3}{125}\]

OpenStudy (anonymous):

Still no? How about this... \[\huge \frac{8a^3}{125}=\frac{2^3a^3}{5^3}\] ???

OpenStudy (anonymous):

(8a^9)/(125a^6) = (8/125)(a^(9-6)) = (8/125)a^3

OpenStudy (anonymous):

Peter I had that it just seemed to easy of an answer.

OpenStudy (anonymous):

Well, awesome, then :) Always believe in yourself!!! :D

OpenStudy (anonymous):

Here's another. Write 32 x^5y^5 with only one exponent. What would you get there. I thought (2xy)^5

OpenStudy (anonymous):

And you're right :)

OpenStudy (anonymous):

ANSWER ( 2xy)^5 So yes